A combustion process of a 100 mol of solid mineral with 325 mole of pure oxygen is

producing ferric oxide and sulfur trioxide gases. The solid mineral contains of 82 mole% iron

pyrites (FeS2) and 18 mole% inerts.

a) Write the chemical equation for the combustion process.

b) Calculate;

i.

The limiting reactant.

ii.

The percentage of excess.

iii.

The extent of reaction, ξ if complete reaction occurred


1
Expert's answer
2021-12-21T13:12:04-0500

2FeS2 + 7O2 = 2FeO + 4SO3

1) According to the equation, 2 moles of FeS2 require 7 moles of O2. Therefore,

82 moles of iron pyrites (FeS2) require: n (O2) = 7 x n (FeS2) / 2 = 287 moles.

According to this, FeS2 is the limiting reactant, O2 is in excess.


2) n (O2)excess = 325 - 287 = 38 moles

n (O2)%excess = 38 / 287 x 100 = 13.2%


3) Extent of reaction is a quantity that measures the extent to which the reaction proceeds.